Which of the following represents the integral  "\displaystyle{\int \frac{x}{\sqrt{1+x^2}}dx}"

1) "\dfrac{x^2}{\sqrt{1+x}}+C"

2) "\ln{(1+x^2)}+C"

3) "\sqrt{1+x^2}+C"

4) "\dfrac{x^2}{\sqrt{1+x^2}}+C"


1
Expert's answer
2021-08-23T18:03:02-0400

We rewrite the integral as: "\int\frac{x}{\sqrt{1+x^2}}dx=\frac{1}{2}\int\frac{d(1+x^2)}{\sqrt{1+x^2}}" and make the change of variables: "z=1+x^2". We receive: "\int\frac{x}{\sqrt{1+x^2}}dx=\frac{1}{2}\int\frac{d(1+x^2)}{\sqrt{1+x^2}}=\frac{1}{2}\int\frac{dz}{\sqrt{z}}"

We integrate the latter and receive: "\frac{1}{2}\int\frac{dz}{\sqrt{z}}=\sqrt{z}+C=\sqrt{1+x^2}+C,C\in{\mathbb{R}}".